# Quantum annealing versus classical solvers: Applications, challenges and limitations for optimisation problems

### v2

@@ 1 Introduction
The field of Quantum Computing (QC) has seen great improvements over the last decade with companies like IBM and Google investing heavily in order to develop resources capable of solving challenging computational problems [1, 2]. Although rapid progress coincides with modern times, the idea of using “simple” systems governed by Quantum Mechanics (QM) to perform computation was proposed in the 1980s by both Richard Feynman [3] and Paul Benioff [4] independently. There are numerous ways to approach building a quantum computer, but the two main categories are gate-based and annealing-based QC. Just like bits in a classical computer, quantum computers use quantum bits, commonly referred to as qubits. Over the years, many different technologies have been suggested and tested as physical implementations of qubits. The first attempts in the 90s used nuclear magnetic resonance [5], later, a plethora of different ideas has shown promise such as cold atoms [6], trapped ions [7], photon-based [8] and others [9, 10]. The current largest quantum annealers are pioneered by D-Wave [11] and consist of loops of superconducting material [12]. It has been shown that in specific cases, QC can be exponentially faster than Classical Computing (CC), such as Shor’s algorithm [13], or the quantum Fourier transform algorithm [14].
The fundamental idea behind QC is to use quantum phenomena such as superposition and entanglement to solve computationally difficult problems. In addition to solving computationally expensive problems, QC may be suited for solving problems that classical computers cannot, such as simulating nature [15, 16], as these systems are governed by the same rules of QM. It has been shown that for several NP-hard problems, where classical approaches fail to efficiently find a solution, quantum computing can be a promising way to solve them [17, 18, 19, 20].
@@
Although quantum annealers are easier to scale towards more qubits, they are far more limited in their applications due to the lack of controllability of the individual qubits. They are, however, especially well suited for optimisation problems where scalability is one of the most determining factors to solve relevant, modern, real-world problems. For a more thorough overview of the synergies between operational research and quantum computing, see [22, 23].
@@
Although D-Wave’s quantum annealers have been applied to solve problems such as rescheduling of urban railroads [31], traffic flow optimisation [32], and dynamic portfolio optimisation [33], the performance or capability of quantum annealers in solving different classes of optimisation problems compared with industrial leading classical solvers is not clear [34]. As pointed out in [31], the current implementations of quantum annealing in optimisation are limited in size and not yet upscaled to real-world situations. Furthermore, the exploration of quantum annealing in scheduling problems is limited. Moreover, given the rapid evolution of QC, regular benchmarking becomes crucial to highlight emerging developments and showcase the current state of this highly promising field, as emphasised by [23]. This research aims to address this gap by utilizing the largest publicly available adiabatic quantum annealers, last updated in August 2023, to solve various optimization problems, including integer/binary linear programming (BLP) with linear and quadratic constraints, integer/binary quadratic programming (BQP), and mixed-integer linear programming. To assess performance, we systematically compare D-Wave’s Advantage quantum annealers, as described in [35], with state-of-the-art classical solvers, namely CPLEX [36], Gurobi [37], and IPOPT [38]. Recognizing the limitations of fully quantum computing algorithms, we employ D-Wave’s hybrid solvers [39]. This approach seamlessly integrates the strengths of both classical and quantum computation, as explained in more detail in Section 2.5. This study serves as a valuable resource for researchers across various fields dealing with optimization problems, enabling them to assess the suitability of the current D-Wave’s quantum annealers for their respective applications.
The subsequent sections of the paper are organized as follows: Section 2 provides a brief introduction to the physical and technical background of quantum annealing, encompassing the adiabatic theorem, and D-Wave’s architecture in their treatment of constraints and the hybrid computing approach employed. Section 3 presents various case studies for different classes of optimization problems. Then, Section 4 showcases the application of D-Wave’s hybrid constrained quadratic model (CQM) solver [40] on an energy system problem formulated as Mixed-Integer Linear Programming (MILP).
@@ 2 Quantum annealing
Simulated Annealing (SA) is used for approximating the global optimal solution, often used when an approximate global solution is preferred over an exact local minimum [41]. The fundamental principle of SA involves a two-phase process: initially “heating” the system, followed by a gradual “cooling” that guides it into its globally lowest energy state. It’s important to emphasize that this “heating” and “cooling” are metaphorical concepts, as no physical heating or cooling is applied; rather, the entire process is mathematically simulated. The inspiration for SA comes from the physical annealing process often used to change the characteristics of different materials [42]. D-Wave utilizes a quantum variation of this process, known as quantum annealing. In QA, the quantum mechanical phenomenon of tunnelling through a potential barrier, rather than the classical process of “climbing” over it, is utilized to escape local minima and potentially find a global one, as explained by Kadowaki and Nishimori (1998) [43]. This fundamental characteristic is pivotal in understanding the potential computational advantages of quantum annealers over their classical counterparts, particularly because the “hill climbing” process in SA can be time-consuming for a substantial subset of problems.
@@
QUBO problems are very closely related to the Ising model (Equation 2) and can be easily transformed into its form by applying an affine transformation $x\mapsto 2x-1$, effectively mapping $x\in\{0,1\}$ to $\{-1,1\}$. With this transformation, the variables $x_{i},x_{j}$ can be represented as the classical spin variables $s_{i}=\pm 1$. So the problem now describes a classical Ising spin system
@@
with spin interaction $J_{i,j}$ and external field $h_{i}$. Traditionally, the Ising model is used in the field of statistical mechanics where the variables represent the binary states of a particle that can either be “spin down” or “spin up”. In quantum mechanics, these variables are replaced with the Pauli operators. By performing this operation, $s$ is mapped to $\frac{\sigma_{z}}{2}$, and the projection of this spin state onto some well-defined axis is considered instead. These spin-projecting operators do not necessarily commute and thus, for some combinations of these operators, unitarity is lost and the process is no longer reversible. Such a system evolves according to the laws of quantum mechanics and the final state eigenvalues correspond to the energy states of the classical Ising model. The Hamiltonian can then be written as
@@
where $B(s)$ is a prefactor with dimensions of energy and $J$ is a coupling coefficient determining the relative strength of the nearest neighbours to the external field. Here, $\sigma_{z}$ is the Pauli matrix with respect to an external magnetic field in the z-direction. D-Wave made use of this close relation and built their machines as a physical realization of such a quantum Ising spin system in a transverse field. However, for quantum computations with such a system to be possible, the machines have to be designed in a way that they can make use of the adiabatic theorem [21], where Equation 3 needs some initial Hamiltonian as a starting point.
@@ 2.1 Adiabatic theorem
In quantum mechanics, operators are used to extract information about physical systems. The most relevant of these operators in our case is the Hamilton operator $H$, usually just referred to as the Hamiltonian. The Hamilton corresponds to the total energy of the quantum mechanical system. By assuming that the time-dependent part varies sufficiently slow and that there is a big enough gap between the eigenstates of $H$, the solutions to the time-dependent Schödinger equation [44] are obtained in terms of the eigenfunctions of the instantaneous Hamiltonian,
@@
where $E_{a}$ is the energy of eigenstate $\psi_{a}$. The scale of “sufficiently slow” depends on the specific energy gap between the quantum states of the system and can range from less than a microsecond to a few seconds. If $H(t)$ varies slowly in time, a system initially in a non-degenerate state $\psi(t=t_{0})$ with energy $E_{a}(t=t_{0})$ will evolve into the corresponding state $\psi_{a}(t)$ with energy $E_{a}(t)$ at a later time $t$ without making any transitions between energy levels. This is known as the adiabatic theorem and is the fundamental principle utilized in quantum annealing. For a more detailed description of the adiabatic theorem and a complete proof, see [21].
To perform calculations using a quantum annealer, the physical collection of qubits, hereafter referred to as the system, is prepared in an easy-to-solve initial Hamiltonian $H_{I}$. After the system has been prepared in $H_{I}$, external magnetic fields are applied to change the initial $H_{I}$ into a complex Hamiltonian $H_{F}$. Suppose this process follows the adiabatic theorem described above. In that case, the system will remain in its ground state for the duration of the anneal and the final state will be the solution of the complex Hamiltonian. The procedure can slowly anneal the system from the quantum Hamiltonian (Equation 5) to the solution of the classical Ising spin system shown in Equation 2. After the annealing process, the collapsed spin states will present a low-energy solution. The annealing quantum Ising Hamiltonian is as follows:
@@
The annealing process begins at $t=0$ and ends at time $t=t_{f}$, where $A(0)\gg B(0)$ and $A(t_{f})\ll B(t_{f})$. At time $t=t_{f}$, the qubits have been dephased to a classical system and the Pauli matrix can be replaced by the classical $s_{i}=\pm 1$, resulting in Equation 2.
@@ 2.2 D-Wave architecture
The most powerful QPU available is the Pegasus topology [46]. In the Pegasus topology, each qubit is coupled to 15 other qubits, which is a great improvement from the 6 couplings of the previous generation. It consists of a square lattice of unit cells which themselves contain 24 qubits. The Advantage QPU, used for the computations performed in this work, is a 16x16 lattice of unit cells, totalling $24\times 16\times(16-1)=5760$ qubits. Due to the densely packed spatial arrangement of qubits and the physical manufacturing process, it is expected that some qubits may exhibit suboptimal performance. D-Wave estimates that approximately 5% of the qubits in the Pegasus topology may have imperfections. Therefore, it is crucial to execute the annealing process repeatedly in succession to mitigate the noise introduced by these faulty qubits. This dense graph structure of the Pegasus systems has allowed for a significant increase in the number of useful qubits, making the Advantage QPU the most promising system to solve the optimisation problems presented in Sections 3.1, 3.4 and 3.3. Due to their physical implementation, quantum annealers can only encode BLP and BQP problems, so terms and constraints have to be transformed into this form [34]. For detailed information on Pegasus and other D-Wave topologies refer to[47]. Additionally, the next generation of topology called Zephyr [48] has already been announced with even higher connectivity. When working with any of these topologies, the embedding of the problem onto the physical hardware is of high importance. While D-Wave does provide a large amount of functionalities to help in accomplishing this process, this is no easy task and requires significant knowledge of the architecture and the problem at hand.
@@ 2.3 Constraints
where $b$ is some constant that has to be met. However, when confronted with a linear inequality constraint of the form $P({\bf{x}})=\left(\sum_{i}^{N}a_{i}x_{i}-b\right)\leq 0$, auxiliary slack variables, ${\bf{w}}=\{w_{j}:\forall j\in J\}$ are used to transform inequality into equality constraints [50] so that Equation 6 can be written as
@@
The amount of slack variables $W$ needed can be determined based on the coefficients $a_{i}$ and $b_{i}$, with an upper limit of $W=b$.
@@ 2.4 State of the art of quantum annealing
The most capable quantum annealers as of today are owned by D-Wave. Therefore, in this paper, we focus on analysing and comparing the performance of their most advanced system against that of classical computers. While D-Wave promotes the capabilities of their machines in solving optimisation problems with 1 million variables and 100,000 constraints [35], it is essential to note that these claims are strongly problem-dependent and not currently realizable simultaneously. The limit of 1 million variables is only valid for unconstrained problems with the possibility of adding constraints via penalties manually. For constrained models, the Advantage system can handle 500,000 variables and 100,000 constraints. These numbers, however, are only the case for D-Wave’s hybrid solvers, which split the problem into different partitions, where some are solved classically and others on the QPU depending on how suited these subproblems are for either of the systems. As an example, continuous variables are usually run on classical computers, as here quantum annealers are not suspected to provide any computational advantage [53]. For pure quantum computation, without any partitioning, the capabilities of the Advantage system are significantly reduced and dependent on the exact embedding of the problem onto the QPU topology. Additionally, the larger the problem size, the “smaller” the logical qubits get, resulting in reduced connectivity and poorer solution quality. The maximum number of variables for a fully connected QUBO on the Advantage hardware is currently at 180, meaning the architecture can only support up to this amount (not considering any faulty qubits) of fully interacting logical variables [54]. For larger fully connected QUBOs, D-Wave uses hybrid workflows and partitioning of the problem into several sub-problems.

### v3

# Quantum annealing applications, challenges and limitations for optimisation problems compared to classical solvers

@@ 1 Introduction
The field of Quantum Computing (QC) has seen great improvements over the last decade with companies like IBM and Google investing heavily in order to develop resources capable of solving challenging computational problems [1, 2]. Recently, Microsoft also announced their new topological qubit design [3], showcasing a new path in obtaining more error-resistant, scalable quantum computers. Although rapid progress coincides with modern times, the idea of using “simple” systems governed by Quantum Mechanics (QM) to perform computation was proposed already in the 1980s by both Richard Feynman [4] and Paul Benioff [5] independently. There are numerous ways to approach building a quantum computer, but the two main categories are gate-based and annealing-based QC. Just like bits in a classical computer, quantum computers use quantum bits, commonly referred to as qubits. Over the years, many different technologies have been suggested and tested as physical implementations of qubits. The first attempts in the 90s used nuclear magnetic resonance [6], later, a plethora of different ideas has shown promise such as cold atoms [7], trapped ions [8], photon-based [9] and others [10, 11]. The current largest quantum annealers are pioneered by D-Wave [12] and consist of loops of superconducting material [13]. It has been shown that in specific cases, QC can be exponentially faster than Classical Computing (CC), such as for Shor’s algorithm [14], or the quantum Fourier transform algorithm [15].
The fundamental idea behind QC is to use quantum phenomena such as superposition and entanglement to solve computationally difficult problems. In addition to solving computationally expensive problems, QC may be suited for solving problems that classical computers cannot, such as simulating nature [16, 17], as these systems are governed by the same rules of QM. It has been shown that for several NP-hard problems, where classical approaches fail to efficiently find a solution, QC can be a promising way to solve them [18, 19, 20, 21].
@@
Although quantum annealers are easier to scale towards more qubits, they are far more limited in their applications due to the lack of controllability of the individual qubits. While in gate-based quantum computers, qubits can be controlled and manipulated with precise unitary operations, the annealing-based computers of D-Wave controls theirs with tunable biases and the coupling strengths between them. D-Wave’s quantum annealers are not universal quantum computers, and therefore can not run arbitrary algorithms. They are, however, especially well suited for optimisation problems where scalability is one of the most determining factors in solving relevant, modern, real-world problems. For a more thorough overview of the synergies between operational research and quantum computing, see [23, 24].
@@
Although D-Wave’s quantum annealers have been applied to solve problems such as rescheduling of urban railroads [32], traffic flow optimisation [33], and dynamic portfolio optimisation [34], the performance or capability of quantum annealers in solving different classes of optimisation problems compared with industrial leading classical solvers is not clear [35]. As pointed out in [32], the current implementations of quantum annealing in optimisation are limited in size and not yet upscaled to real-world situations. Furthermore, the exploration of quantum annealing in scheduling problems is limited. Given the rapid evolution of QC, regular benchmarking becomes crucial to highlight emerging developments and showcase the current state of this highly promising field, as emphasised by [24]. Benchmarks like [36] have shown that D-Wave can compete with classical solvers like Tabu Search [37] and Simulated Annealing (SA) [38] for energy coalition formation and is ahead of gate-based applications like the Quantum Approximate Optimisation Algorithm [39]. However, no benchmark of industry-leading solvers has been conducted that evaluates their performance against that of D-Wave’s hybrid solvers for different optimisation problem categories.
This research aims to address this gap by utilising the largest publicly available adiabatic quantum annealers, last updated in August 2023, to solve various optimisation problems, including integer/binary linear programming (BLP) with linear and quadratic constraints, integer/binary quadratic programming (BQP), and mixed-integer linear programming. To assess performance, we systematically compare D-Wave’s Advantage quantum annealers, as described in [40], with the state-of-the-art classical solvers CPLEX [41], Gurobi [42], and IPOPT [43]. Recognising the limitations of fully QC algorithms, we employ D-Wave’s hybrid solvers [44]. This approach seamlessly integrates the strengths of both classical and quantum computation, as explained in more detail in Section 2.5. This study serves as a valuable resource for researchers across various fields dealing with optimisation problems, enabling them to assess the suitability of current D-Wave quantum annealers for their respective applications.
The subsequent sections of the paper are organised as follows: Section 2 provides a brief introduction to the physical and technical background of quantum annealing, encompassing the adiabatic theorem and D-Wave’s architecture in their treatment of constraints and the hybrid computing approach employed. Section 3 presents various case studies for different classes of optimisation problems. Then, Section 4 showcases the application of D-Wave’s hybrid constrained quadratic model (CQM) solver [45] on an energy system problem formulated as Mixed-Integer Linear Programming (MILP).
@@ 2 Quantum annealing
Simulated Annealing (SA) is used for approximating the global optimal solution, often used when an approximate global solution is preferred over an exact local minimum [38]. The fundamental principle of SA involves a two-phase process: initially “heating” the system, followed by a gradual “cooling” that guides it into its globally lowest energy state. It is important to emphasise that this “heating” and “cooling” are metaphorical concepts, as no physical heating or cooling is applied; rather, the entire process is mathematically simulated. The inspiration for SA comes from the physical annealing process often used to change the characteristics of different materials [46]. D-Wave utilises a quantum variation of this process, known as quantum annealing. In quantum annealing, the quantum mechanical phenomenon of tunnelling through a potential barrier, rather than the classical process of “climbing” over it, is utilised to escape local minima and potentially find a global one, as explained by Kadowaki and Nishimori (1998) [47]. This fundamental characteristic is pivotal in understanding the potential computational advantages of quantum annealers over their classical counterparts, particularly because the “hill climbing” process in SA can be time-consuming for a substantial subset of problems.
@@
QUBO problems are very closely related to the Ising model (Equation 2) and can be easily transformed into this form by applying an affine transformation $x\mapsto 2x-1$, effectively mapping $x\in\{0,1\}$ to $\{-1,1\}$. Under this transformation, the variables $x_{i},x_{j}$ correspond to classical spin variables $s_{i}=\pm 1$ leading to the classical Ising spin system:
@@
where $J_{i,j}$ represents the spin interaction strength and $h_{i}$ external field. Traditionally, the Ising model is used in the field of statistical mechanics where the variables represent the binary states of a particle that can either be “spin down” or “spin up”.
In quantum mechanics, these variables are replaced by Pauli operators $\sigma_{x,y,z}$. Specifically, spin states are often expressed by terms of the Pauli $\sigma_{z}$ matrix which correspond to spin projections along the z-axis. The Hamiltonian describing such a quantum system takes the form:
@@
where $B(s)$ is a prefactor with dimensions of energy and $J_{i,j}$ determines the interaction strength between neighboring spins. Since all the terms in this Hamiltonian involve $\sigma_{z}$ operators, they mutually commute. However, quantum annealers such as D-Wave introduce an additional transverse field involving $\sigma_{x}$ terms, which do not necessarily commute with $\sigma_{z}$. This non-commutativity is crucial for quantum annealing to leverage the adiabatic theorem [22], requiring an initial Hamiltonian with a dominant transverse field component as a starting point before gradually transitioning into the problem Hamiltonian.
@@ 2.1 Adiabatic theorem
In quantum mechanics, operators are used to extract information about physical systems. The most relevant of these operators in our case is the Hamilton operator $H$, usually just referred to as the Hamiltonian. The Hamiltonian corresponds to the total energy of the quantum mechanical system. By assuming that the time-dependent part varies sufficiently slowly and that there is a big enough gap between the eigenstates of $H$, the solutions to the time-dependent Schödinger equation [48] are obtained in terms of the eigenfunctions of the instantaneous Hamiltonian,
@@
where $E_{a}$ is the energy of eigenstate $\psi_{a}$. The scale of “sufficiently slow” depends on the specific energy gap between the quantum states of the system and can range from less than a microsecond to a few seconds. If $H(t)$ varies slowly in time, a system initially in a non-degenerate state $\psi(t=t_{0})$ with energy $E_{a}(t=t_{0})$ will evolve into the corresponding state $\psi_{a}(t)$ with energy $E_{a}(t)$ at a later time $t$ without making any transitions between energy levels. This is known as the adiabatic theorem and is the fundamental principle utilised in quantum annealing. For a more detailed description of the adiabatic theorem and a complete proof thereof, see [22].
To perform calculations using a quantum annealer, the physical collection of qubits, hereafter referred to as the system, is prepared in an easy-to-solve initial Hamiltonian $H_{I}$. After the system has been prepared in $H_{I}$, external magnetic fields are applied to change the initial $H_{I}$ into a complex Hamiltonian $H_{F}$. Suppose this process follows the adiabatic theorem described above. In that case, the system will remain in its ground state for the duration of the anneal, and the final state will be the solution of the complex Hamiltonian. As $[\sigma_{x},\sigma_{z}]\neq 0$, $H_{I}$ and $H_{F}$ do not share a common set of eigenstates. As a result, the system must undergo quantum transitions to evolve from one eigenstate to another. If the two Hamiltonians did commute, the eigenstates would remain unchanged throughout the evolution, preventing any quantum dynamics from occurring. The procedure can slowly anneal the system from the quantum Hamiltonian (Equation 5) to the solution of the classical Ising spin system shown in Equation 2. After the annealing process, the collapsed spin states will present a low-energy solution. The annealing quantum Ising Hamiltonian is as follows:
@@
The annealing process begins at $t=0$ and ends at time $t=t_{f}$, where $A(0)\gg B(0)$ and $A(t_{f})\ll B(t_{f})$. At time $t=t_{f}$, the qubits have been de-phased to a classical system and the Pauli matrix can be replaced by the classical $s_{i}=\pm 1$, resulting in Equation 2.
@@ 2.2 D-Wave architecture
The most powerful QPU available is the Pegasus topology [50]. In the Pegasus topology, each qubit is coupled to 15 other qubits, which is a great improvement from the 6 couplings of the previous generation. It consists of a square lattice of unit cells which themselves contain 24 qubits. The Advantage QPU, used for the computations performed in this work, is a 16x16 lattice of unit cells, totalling $24\times 16\times(16-1)=5760$ qubits. Due to the densely packed spatial arrangement of qubits and the physical manufacturing process, it is expected that some qubits may exhibit suboptimal performance. D-Wave estimates that approximately 5% of the qubits in the Pegasus topology may have imperfections. This dense graph structure of the Pegasus systems has allowed for a significant increase in the number of useful qubits, making the Advantage QPU the most promising system to solve the optimisation problems presented in Sections 3.1, 3.4 and 3.3.
Quantum devices like those from D-Wave are subject to several causes of errors, including thermal noise, control errors, flux bias drifts and decoherence, which can lead to wrong solutions. To mitigate some of these, D-Wave has implemented Drift Correction techniques [51] to compensate for slow fluctuations in qubit parameters by periodically measuring and adjusting flux biases to maintain computational accuracy. Additionally, D-Wave has demonstrated on the new Advantage-2 prototype Zero-Noise Extrapolation (ZNE) [52] using controlled noise amplification and extrapolation to estimate error-free results. However, ZNE has not yet been implemented on commercially available systems. Generally, D-Wave uses statistical sampling by running multiple annealing cycles to generate a distribution of solutions, mitigating the noise introduced by these faulty qubits.
Due to their physical implementation, current quantum annealers can only encode BLP and BQP problems, so terms and constraints have to be transformed into this form [35]. For detailed information on Pegasus and other D-Wave topologies, refer to[53]. Additionally, the next generation of topology, called Zephyr [54], has already been announced with even higher connectivity. When working with any of these topologies, the embedding of the problem onto the physical hardware is of high importance. While D-Wave does provide a large amount of functionalities to help in accomplishing this process, this is no easy task and requires significant knowledge of the architecture and the problem at hand.
@@ 2.3 Constraints
where $b$ is some constant that has to be met. However, when confronted with a linear inequality constraint of the form $P({\bf{x}})=\left(\sum_{i}^{N}a_{i}x_{i}-b\right)\leq 0$, auxiliary slack variables, ${\bf{y}}=\{y_{j}:\forall j\in J\}$ and their corresponding weights $w_{j}$ are used to transform inequality constraints into equality constraints [56] so that Equation 6 can be written as
@@
The number of slack variables $W$ needed can be determined based on the coefficients $a_{i}$ and $b_{i}$, with an upper limit of $W=b$.
@@ 2.4 State of the art of quantum annealing
The most capable quantum annealers as of today are owned by D-Wave. Therefore, in this paper, we focus on analysing and comparing the performance of their most advanced system against that of classical computers. While D-Wave promotes the capabilities of their machines in solving optimisation problems with 1 million variables and 100,000 constraints [40], it is essential to note that these claims are strongly problem-dependent and not currently realizable simultaneously. The limit of 1 million variables is only valid for unconstrained problems with the possibility of adding constraints via penalties manually. For constrained models, the Advantage system can handle 500,000 variables and 100,000 constraints. These numbers, however, are only the case for D-Wave’s hybrid solvers, which split the problem into different partitions, where some are solved classically and others on the QPU depending on how suited these sub-problems are for either of the systems. As an example, continuous variables are usually run on classical computers, as here quantum annealers are not suspected to provide any computational advantage [59]. For pure quantum computation, without any partitioning, the capabilities of the Advantage system are significantly reduced and dependent on the exact embedding of the problem onto the QPU topology. Additionally, the larger the problem size, the “smaller” the logical qubits get, resulting in reduced connectivity and poorer solution quality. The maximum number of variables for a fully connected QUBO on the Advantage hardware is currently at 180, meaning the architecture can only support up to this amount (not considering any faulty qubits) of fully interacting logical variables [60]. For larger fully connected QUBOs, D-Wave uses hybrid workflows and partitioning of the problem into several sub-problems.